# NJL-Chiral Soliton and the Nucleon Equation of State at supra-saturation density: Impact of Chiral Symmetry Restoration

### v1

@@ II.3 Numerical Implementations
Before going into the vast connection between nucleon properties of the NJL model parameters, we establish the reliability of our numerical framework for the NJL-soliton model, validating our implementation by reproducing the results of Ref. [58]. The skyrmion framework, as detailed in Ref. [58], provides soliton solutions in vacuum. In our NJL soliton framework, this is equivalently characterized by the scalar field ${\cal S}=M_{0}$. In this section, we adopt the below-mentioned meson parameters to recover the skyrmion framework from Ref. [58] to ensure consistency and direct comparison to their findings. Specifically, we use the following meson parameters to obtain the soliton solutions at vacuum following the methodology given in section II.1:
@@
Furthermore, for the vector and iso-vector mesons, we use $m_{\omega}=m_{\rho}=783$ MeV and $g=6$. These substitutions align the NJL-soliton equations, see eqs. 18, 23 and 24, with the expressions derived in Ref. [58, 57], enabling a direct comparison.
In the relaxation method, we first employ a compactified coordinate transformation, $t=\tanh(r)\in[0,1]$ ( where $r$ is supposed to be given in fm), to handle the boundary condition at $r\to\infty$ and speed up the convergence. This transformation ensures numerical stability, despite the rapid variation of the tanh function, as the mapping allows us to reach $r_{\infty}\sim 10$ fm for $1-t\sim 10^{-9}$, which sufficiently captures the soliton tail, given that the characteristic nucleon radius is $\sim 1$ fm. A second-order central finite differences approximation is considered for the derivatives in the discretized coordinate system. The differential eqs. 23, 24 and 25 were reformulated as a nonlinear algebraic system, and a residual vector was constructed by subtracting the right-hand sides of the field equations eqs. 23, 24 and 25 from the finite-difference approximations. The resulting nonlinear algebraic system is solved iteratively by minimizing the residual vector constructed from the discretized field equations. For initial guesses, we consider physically motivated profile functions inspired by prior literature. Exponentially decaying forms such as $F(r)=\pi e^{-M_{\pi}r}$, $G(r)=-2e^{-M_{v}r}$, and $\omega(r)=\omega_{0}e^{-M_{v}r}$ yield reasonable convergence. However, we observe faster and more stable convergence with tanh-type trial functions: $F(r)=\pi(1-\tanh(M_{\pi}r))$, $G(r)=-2(1-\tanh(M_{v}r))$, and $\omega(r)=\omega_{0}(1-\tanh(m_{\omega}r))$. These satisfy the boundary conditions and exhibit the correct asymptotic behavior implied by eqs. 23 and 25.
To compute the soliton’s physical properties in the radial domain with minimal errors, we used the converged solution in the compactified coordinate $t=\tanh(r)$ as an initial guess for the relaxation method in $r$. Because we begin with a smooth, converged solution in $t$, convergence in $r$ is both stable and precise resulting as a polishing algorithm. A thorough numerical accuracy and error analysis is presented in Appendix A. The meson field profiles, obtained using both the shooting and relaxation methods, are presented in fig. 1. As shown in fig. 1, the field profiles obtained with the relaxation method agree with those obtained from the shooting method within $\leq 5\%$ over the relevant soliton radius range. The excellent agreement between these results confirms the robustness of our implementation.
@@
Using the parameters from Ref. [58], we computed the normalized baryon and isoscalar charge density profiles, shown in fig. 2. The baryon density $\rho_{B}(r)$ reflects the valence quark contributions, defining the hard-core radius $r_{B}$, while the isoscalar density $\rho_{S}(r)$ captures the sea quark contributions, characterizing the isoscalar charge radius $r_{S}$.
Figure 2: Baryon ($\rho_{B}(r)$) and isoscalar ($\rho_{S}(r)$) charge densities, scaled by $4\pi r^{2}$, as functions of $r$. The radial axis is limited to $r_{\rm max}=1.5$ fm for visibility and comparison with Ref. [58]. For clarity and comparison with Ref. [58], the isoscalar charge density is shown as $\rho_{S}(r)=-2gF_{\pi}^{2}\omega(r)$.
@@
The baryon charge radius, $\sqrt{\langle r^{2}_{B}\rangle}\approx 0.49$ fm, agrees well with Ref. [58]. However, a discrepancy arises in the isoscalar charge radius, where we obtain $\sqrt{\langle r^{2}_{S}\rangle}\approx 0.80$ fm, compared to $1.03$ fm reported in Ref. [58]. This difference, approximately 22%, can be reconciled using the relation among $\langle r^{2}_{B}\rangle$ and $\langle r^{2}_{S}\rangle$ within the form factor argument as given in Refs. [57, 58, 53]:
@@
with $m_{\omega}=783$ MeV, we find $\langle r^{2}_{S}\rangle\approx 0.63~\mathrm{fm}^{2}$, or $\sqrt{\langle r^{2}_{S}\rangle}\approx 0.8$ fm consistent with our result and with the earlier findings in Ref. [57] (see section 4.4 therein) which is the foundation of the model used in Ref. [58].
The energy density $\varepsilon_{\rm{slt}}(r)$ and pressure $p_{\rm{slt}}(r)$ profiles of the soliton are shown in fig. 3. The pressure exhibits the characteristic feature of a positive core region and a negative surface region, ensuring global mechanical stability through the Virial theorem, $\int d^{3}r\,p_{\rm{slt}}(r)=0$. We find the soliton mass $M_{\rm{slt}}=\int\varepsilon_{\rm{slt}}(r)d^{3}r\approx 1452$ MeV, which agrees with the value $\sim 1460$ MeV reported in Ref. [58] within 0.5%, confirming the accuracy of our numerical approach. An interesting quantity is the central ($r=0$) energy density and pressure, hereafter noted $\varepsilon^{c}_{\rm{slt}}$ and pressure $p^{c}_{\rm{slt}}$, which represent the point for which the energy density and the pressure are maximum as well as the point which is the further one from the surface. It is therefore the point that is the least contaminated by surface contributions.
Figure 3: Upper panel shows the scaled (with $4\pi r^{2}$) energy density $\varepsilon_{\rm slt}(r)$ and pressure $p_{\rm slt}(r)$ as functions of $r$. The lower panel shows the variation of $\varepsilon_{\rm slt}(r)$ and $p_{\rm slt}(r)$ as a function of r. The displayed quantities correspond to the soliton solution with the meson parameters of Ref. [58].
In the context of the Skyrme model, it is well understood that the static soliton must be properly quantized in spin and isospin to reproduce the nucleon mass in its ground state  [^2] Though the quantization successfully explains the mass difference between nucleon ($N$) and Delta ($\Delta$) baryon. The spin-isospin quantization increases the soliton mass by $3/8I$, II being the moment of inertia. Furthermore, one needs to account for the center of mass correction for obtaining the mass of the nucleon. . However, the quantization procedure is beyond the scope of the present work. Instead, following the prescription of Ref. [58], we introduce a phenomenological scaling parameter $\chi$, defined at vacuum as,
@@
This scaling allows us to match the observed nucleon mass and enables a meaningful comparison between model predictions and physical observables. Accordingly, we rescale the soliton energy density and pressure profiles as follows [58]:
@@
This rescaling procedure also facilitates direct comparison with the Nucleon EoS shown in Figure 8 of Ref. [58]. Based on their qualitative argument, a hard-core radius of $\sim 0.5$ fm suggests that hard deconfinement may set in around $\sim 8\rho_{0}$ (or possibly earlier). Furthermore, within the hard-confinement regime, it is argued that the thermodynamic pressure of bulk matter can be approximated by the mechanical pressure found at the core of the nucleon.
To isolate the surface effects specific to an isolated nucleon, we consider the pressure distribution near the center of the soliton [58]. As a test of our numerical implementation, we present the effective EoS for the nucleon core—obtained by plotting $p_{\rm slt,\chi}(r)$ as a function of $\varepsilon_{\rm slt,\chi}(r)$ after eliminating the radial coordinate $r$. For context and comparison, we also include the NS $\beta$-equilibrated solution [90] of the SLy4 nuclear interaction [91] and the QHC18 model [92] obtained from the CompOSE [93] repository are also shown in fig. 4.
Figure 4: Effective core EoS of the nucleon: pressure $p_{\rm slt,\chi}(r)$ as a function of energy density $\varepsilon_{\rm slt,\chi}(r)$, obtained by eliminating $r$ from the profiles shown in fig. 3. Red circles indicate specific radial positions. For comparison, the solution [90] of the SLy4 nuclear interaction [91] and QHC18 [92] EoS models are also shown.
III Results
After successfully reproducing the key results from Ref. [58] with great accuracy, we now proceed to our objectives of this work: to investigate the sensitivity of vacuum soliton properties to the underlying NJL model parameters and to investigate the effect of chiral symmetry restoration in a self-consistent framework. The former analysis is presented in detail in section III.1. Furthermore, for the latter, we analyze the effects of chiral symmetry restoration on soliton structure and the corresponding EoS, solely by means of the scalar field in section III.2.1 and in dense nuclear matter in  section III.2.2.
III.1 NJL Parameters and Soliton in Vacuum
@@
The parameter set labelled NJL-F is constructed to closely reproduce the vacuum soliton properties reported in Ref. [58], but within the bosonization scheme adopted in this work, consistent with the formulation developed in Ref. [68]. The label “F” in NJL-F refers to “Fukushima et al.”. In addition, we examine two further parameter sets inspired by the work of Ref. [68]. The set NJL-C2 corresponds directly to the parameter set denoted as “NJLSET2” in Ref. [68], while NJL-C1 represents an alternative choice within the same framework. Here, the label “C” stands for “Chanfray”, and the number distinguishes between the different parameter choices. All these parameter sets are summarised in table 1.
Although the NJL-F parametrization reproduces the meson properties (in particular the (iso-)vector sector), reasonably well with the values used in Ref. [58], we find that the alternative parametrization NJL-FKW, listed in table 1, exactly reproduces both the scalar and vector meson properties. Consequently, the soliton EoS obtained with NJL-FKW exactly reproduces the nucleon EoS of Ref. [58], consistent with the results shown in figs. 7(b) and 7(a). Despite this apparent advantage, we do not adopt NJL-FKW in our further analysis because the corresponding cut-off parameter is quite large, $\Lambda>1~\text{GeV}$, along with the current quark mass, rather than being small $m_{q}\sim 1.5$ MeV. Nevertheless, the omission of this special parametrization does not affect the central motivation, main findings, or qualitative conclusions of the present work.
@@
Figure 5: Solid lines in different panels show the meson field profiles for different NJL parameters as noted in the sub-panel. For comparison, the field profiles corresponding to the parameters of Ref. [58] are also shown in dashed curves.
@@
The meson field profiles for the NJL-F, NJL-C1, and NJL-C2 parameter sets are shown in fig. 5, alongside the profiles corresponding to the parameters from Ref. [58] for comparison. Key soliton properties such as soliton mass $M_{\rm slt}$, baryon charge radius $\sqrt{\langle r_{B}^{2}\rangle}$, and isoscalar charge radius $\sqrt{\langle r_{S}^{2}\rangle}$ at vacuum are summarized in table 2. As the vector meson mass $M_{v}$ increases across the parameter sets, the $\omega$-field becomes more localized near the core of the soliton, as evident in fig. 5. This localization leads to a sharper radial decay, resulting in a reduced isoscalar charge radius $\sqrt{\langle r_{S}^{2}\rangle}$, as shown in fig. 6 and quantified in table 2. In contrast, the baryon charge radius $\sqrt{\langle r_{B}^{2}\rangle}$, associated with the valence quark distribution, remains relatively insensitive to changes in $M_{v}$, indicating the robustness of the hard-core structure of the soliton.
@@
We also present the nucleon-scalar coupling $g_{S}$ and the scalar susceptibility $C_{\rm NS}$ within the NJL-soliton framework, defined as in Refs. [67, 68]:
@@
We observe a notably small nucleon-scalar coupling [^3] Alternatively, if one evolve the soliton only through the scalar sector of the NJL-soliton model ($F_{\pi}({\cal S}),M_{\pi}({\cal S})$), while keeping the vector parameters fixed at their vacuum values $M_{v}({\cal S})=M_{v}(M_{0}),\ g_{v}({\cal S})=g_{v}(M_{0})$], as done in several works, see Refs. [77, 78], leads to a different nucleon scalar coupling, as discussed in appendix B. constant within the present framework, accompanied by a negative scalar susceptibility. The appearance of a negative susceptibility is consistent within the NJL-Soliton/skyrmion model. However, we anticipate that incorporating confinement effects, for instance within the chiral confining model, will mitigate this behavior, potentially yielding a positive susceptibility more in line with the results reported in Refs. [67, 68].
@@
Although we adopt a normalization procedure involving the energy density, specifically multiplying by the factor $\chi$, to reproduce the physical nucleon mass at vacuum (as outlined in section II.2), we present and compare the resulting EoS both with and without this approach. The EoSs for the various NJL parameter sets, obtained by eliminating the radial coordinate $r$ from the profiles of energy density $\varepsilon_{\rm slt}(r)$ and pressure $p_{\rm slt}(r)$, are shown in fig. 7(a), alongside the reference EoS corresponding to the exact parameters of Ref. [58]. The impact of the scaling parameter $\chi$ on the EoS is illustrated in fig. 7(b), and the scaling approach changes the stiffness of the EoSs.
III.2 Effects of Chiral Symmetry Restoration
III.2.1 Scalar field dependency

### v2

@@ 2.3 Numerical Implementations
Before going into the vast connection between nucleon properties of the NJL model parameters, we establish the reliability of our numerical framework for the NJL-soliton model, validating our implementation by reproducing the results of Ref. Fukushima2020 . The skyrmion framework, as detailed in Ref. Fukushima2020 , provides soliton solutions in vacuum. In our NJL soliton framework, this is equivalently characterized by the scalar field ${\cal S}=M_{0}$. In this section, we adopt the below-mentioned meson parameters to recover the skyrmion framework from Ref. Fukushima2020 to ensure consistency and direct comparison to their findings. Specifically, we use the following meson parameters to obtain the soliton solutions at vacuum following the methodology given in Section 2.1:
@@
Furthermore, for the vector and iso-vector mesons, we use $m_{\omega}=m_{\rho}=783$ MeV and $g=6$. These substitutions align the NJL-soliton equations, see Eqs. (18), (23), and (24), with the expressions derived in Ref. Fukushima2020 ; Meissner1987 , enabling a direct comparison.
In the relaxation method, we first employ a compactified coordinate transformation, $t=\tanh(r)\in[0,1]$ (where $r$ is expressed in fm), to handle the boundary condition at $r\to\infty$ and speed up the convergence. This transformation ensures numerical stability, despite the rapid variation of the tanh function, as the mapping allows us to reach $r_{\infty}\sim 10$ fm for $1-t\sim 10^{-9}$, which sufficiently captures the soliton tail, given that the characteristic nucleon radius is $\sim 1$ fm. A second-order central finite differences approximation is considered for the derivatives in the discretized coordinate system. The differential Eqs. (23), (24), and (25) were reformulated as a nonlinear algebraic system, and a residual vector was constructed by subtracting the right-hand sides of the field equations Eqs. (23), (24), and (25) from the finite-difference approximations. The resulting nonlinear algebraic system is solved iteratively by minimizing the residual vector constructed from the discretized field equations. For initial guesses, we consider physically motivated profile functions inspired by prior literature. Exponentially decaying forms such as $F(r)=\pi e^{-M_{\pi}r}$, $G(r)=-2e^{-M_{v}r}$, and $\omega(r)=\omega_{0}e^{-M_{v}r}$ yield reasonable convergence. However, we observe faster and more stable convergence with tanh-type trial functions: $F(r)=\pi(1-\tanh(M_{\pi}r))$, $G(r)=-2(1-\tanh(M_{v}r))$, and $\omega(r)=\omega_{0}(1-\tanh(m_{\omega}r))$. These satisfy the boundary conditions and exhibit the correct asymptotic behavior implied by Eqs. (23), (25).
To compute the soliton’s physical properties in the radial domain with minimal errors, we used the converged solution in the compactified coordinate $t=\tanh(r)$ as an initial guess for the relaxation method in $r$. Because we begin with a smooth, converged solution in $t$, convergence in $r$ is both stable and precise resulting as a polishing algorithm. A thorough numerical accuracy and error analysis is presented in  A. The meson field profiles, obtained using both the shooting and relaxation methods, are presented in Fig. 1. As shown in Fig. 1, the field profiles obtained with the relaxation method agree with those obtained from the shooting method within $\leq 5\%$ over the relevant soliton radius range. The excellent agreement between these results confirms the robustness of our implementation.
@@
Using the parameters from Ref. Fukushima2020 , we computed the normalized baryon and isoscalar charge density profiles, shown in Fig. 2. The baryon density $\rho_{B}(r)$ reflects the valence quark contributions, defining the hard-core radius $r_{B}$, while the isoscalar density $\rho_{S}(r)$ captures the sea quark contributions, characterizing the isoscalar charge radius $r_{S}$.
Figure 2: Baryon ($\rho_{B}(r)$) and isoscalar ($\rho_{S}(r)$) charge densities, scaled by $4\pi r^{2}$, as functions of $r$. The radial axis is limited to $r_{\rm max}=1.5$ fm for visibility and comparison with Ref. Fukushima2020 . For clarity and comparison with Ref. Fukushima2020 , the isoscalar charge density is shown as $\rho_{S}(r)=-2gF_{\pi}^{2}\omega(r)$.
@@
The baryon charge radius, $\sqrt{\langle r^{2}_{B}\rangle}\approx 0.49$ fm, agrees well with Ref. Fukushima2020 . However, a discrepancy arises in the isoscalar charge radius, where we obtain $\sqrt{\langle r^{2}_{S}\rangle}\approx 0.80$ fm, compared to $1.03$ fm reported in Ref. Fukushima2020 . This difference, approximately 22%, can be reconciled using the relation among $\langle r^{2}_{B}\rangle$ and $\langle r^{2}_{S}\rangle$ within the form factor argument as given in Refs. Meissner1987 ; Fukushima2020 ; ADKINS1984251 :
@@
with $m_{\omega}=783$ MeV, we find $\langle r^{2}_{S}\rangle\approx 0.63~\mathrm{fm}^{2}$, or $\sqrt{\langle r^{2}_{S}\rangle}\approx 0.8$ fm consistent with our result and with the earlier findings in Ref. Meissner1987 (see section 4.4 therein) which is the foundation of the model used in Ref. Fukushima2020 .
The energy density $\varepsilon_{\rm{slt}}(r)$ and pressure $p_{\rm{slt}}(r)$ profiles of the soliton are shown in Fig. 3. The pressure exhibits the characteristic feature of a positive core region and a negative surface region, ensuring global mechanical stability through the Virial theorem, $\int d^{3}r\,p_{\rm{slt}}(r)=0$. We find the soliton mass $M_{\rm{slt}}=\int\varepsilon_{\rm{slt}}(r)d^{3}r\approx 1452$ MeV, which agrees with the value $\sim 1460$ MeV reported in Ref. Fukushima2020 within 0.5%, confirming the accuracy of our numerical approach. An interesting quantity is the central ($r=0$) energy density and pressure, hereafter noted $\varepsilon^{c}_{\rm{slt}}$ and pressure $p^{c}_{\rm{slt}}$, which represent the point for which the energy density and the pressure are maximum as well as the point which is the further one from the surface. It is therefore the point that is the least contaminated by surface contributions.
Figure 3: Upper panel shows the scaled (with $4\pi r^{2}$) energy density $\varepsilon_{\rm slt}(r)$ and pressure $p_{\rm slt}(r)$ as functions of $r$. The lower panel shows the variation of $\varepsilon_{\rm slt}(r)$ and $p_{\rm slt}(r)$ as a function of r. The displayed quantities correspond to the soliton solution with the meson parameters of Ref. Fukushima2020 .
In the context of the Skyrme model, it is well understood that the static soliton must be properly quantized in spin and isospin to reproduce the nucleon mass in its ground state. The static soliton represents a coherent superposition of spin–isospin states and therefore does not correspond to a pure nucleon state. A proper treatment would require collective quantization in spin and isospin, which generates the nucleon–Delta splitting through the moment of inertia and removes spurious rotational contributions  [^2] Though the quantization successfully explains the mass difference between nucleon ($N$) and Delta ($\Delta$) baryon. The spin-isospin quantization increases the soliton mass by $3/8I$, II being the moment of inertia. Furthermore, one needs to account for the center of mass correction for obtaining the mass of the nucleon. . However, our main purpose is to address the in medium effect on the soliton EoS, which is barely discussed in  Fukushima2020 . Furthermore the quantization procedure is beyond the scope of the present work and can be addressed in the future work. Instead, we adopt the same mapping between the soliton model and nucleon as in  Fukushima2020 and introduce a phenomenological scaling parameter $\chi$, defined at vacuum as,
@@
The scaling parameter ($\chi$) should be viewed as a phenomenological correction for spurious collective effects inherent to the soliton, rather than as a substitute for full spin–isospin quantization. This scaling allows us to match the observed nucleon mass and enables a meaningful comparison between model predictions and physical observables. Accordingly, we rescale the soliton energy density and pressure profiles as follows Fukushima2020 :
@@
This rescaling procedure also facilitates direct comparison with the Nucleon EoS shown in Figure 8 of Ref. Fukushima2020 . Based on their qualitative argument, a hard-core radius of $\sim 0.5$ fm suggests that hard deconfinement may set in around $\sim 8\rho_{0}$ (or possibly earlier). Furthermore, within the hard-confinement regime, it is argued that the thermodynamic pressure of bulk matter can be approximated by the mechanical pressure found at the core of the nucleon.
To isolate the surface effects specific to an isolated nucleon, we consider the pressure distribution near the center of the soliton Fukushima2020 . As a test of our numerical implementation, we present the effective EoS for the nucleon core—obtained by plotting $p_{\rm slt,\chi}(r)$ as a function of $\varepsilon_{\rm slt,\chi}(r)$ after eliminating the radial coordinate $r$. For context and comparison, we also include the NS $\beta$-equilibrated solution Gulminelli_2015_SLy4 of the SLy4 nuclear interaction Chabanat:1998 and the QHC18 model Baym_QHC2018 obtained from the CompOSE CompOSE2022 repository are also shown in Fig. 4.
Figure 4: Effective core EoS of the nucleon: pressure $p_{\rm slt,\chi}(r)$ as a function of energy density $\varepsilon_{\rm slt,\chi}(r)$, obtained by eliminating $r$ from the profiles shown in Fig. 3. Red circles indicate specific radial positions. For comparison, the solution Gulminelli_2015_SLy4 of the SLy4 nuclear interaction Chabanat:1998 and QHC18 Baym_QHC2018 EoS models are also shown.
3 Results
After successfully reproducing the key results from Ref. Fukushima2020 with great accuracy, we now proceed to our objectives of this work: to investigate the sensitivity of vacuum soliton properties to the underlying NJL model parameters and to investigate the effect of chiral symmetry restoration in a self-consistent framework. The former analysis is presented in detail in Section 3.1. Furthermore, for the latter, we analyze the effects of chiral symmetry restoration on soliton structure and the corresponding EoS, solely by means of the scalar field in Section 3.2.1 and in dense nuclear matter in Section 3.2.2.
3.1 NJL Parameters and Soliton in Vacuum
@@
The parameter set labelled NJL-F is constructed to closely reproduce the vacuum soliton properties reported in Ref. Fukushima2020 , but within the bosonization scheme adopted in this work, consistent with the formulation developed in Ref. Chanfray2025 . The label “F” in NJL-F refers to “Fukushima et al.”. In addition, we examine two further parameter sets inspired by the work of Ref. Chanfray2025 . The set NJL-C2 corresponds directly to the parameter set denoted as “NJLSET2” in Ref. Chanfray2025 , while NJL-C1 represents an alternative choice within the same framework. Here, the label “C” stands for “Chanfray”, and the number distinguishes between the different parameter choices. All these parameter sets are summarised in Table 1.
Although the NJL-F parametrization reproduces the meson properties (in particular the (iso-)vector sector), reasonably well with the values used in Ref. Fukushima2020 , we find that the alternative parametrization NJL-FKW, listed in Table 1, exactly reproduces both the scalar and vector meson properties. Consequently, the soliton EoS obtained with NJL-FKW exactly reproduces the nucleon EoS of Ref. Fukushima2020 , consistent with the results shown in Figs. 7(b), and  7(a). Despite this apparent advantage, we do not adopt NJL-FKW in our further analysis because the corresponding cut-off parameter is quite large, $\Lambda>1~\text{GeV}$, along with the current quark mass, rather than being small $m_{q}\sim 1.5$ MeV. Nevertheless, the omission of this special parametrization does not affect the central motivation, main findings, or qualitative conclusions of the present work.
@@
Figure 5: Solid lines in different panels show the meson field profiles for different NJL parameters as noted in the sub-panel. For comparison, the field profiles corresponding to the parameters of Ref. Fukushima2020 are also shown in dashed curves.
@@
The meson field profiles for the NJL-F, NJL-C1, and NJL-C2 parameter sets are shown in Fig. 5, alongside the profiles corresponding to the parameters from Ref. Fukushima2020 for comparison. Key soliton properties such as soliton mass $M_{\rm slt}$, baryon charge radius $\sqrt{\langle r_{B}^{2}\rangle}$, and isoscalar charge radius $\sqrt{\langle r_{S}^{2}\rangle}$ at vacuum are summarized in Table 2. As the vector meson mass $M_{v}$ increases across the parameter sets, the $\omega$-field becomes more localized near the core of the soliton, as evident in Fig. 5. This localization leads to a sharper radial decay, resulting in a reduced isoscalar charge radius $\sqrt{\langle r_{S}^{2}\rangle}$, as shown in Fig. 6 and quantified in Table 2. In contrast, the baryon charge radius $\sqrt{\langle r_{B}^{2}\rangle}$, associated with the valence quark distribution, remains relatively insensitive to changes in $M_{v}$, indicating the robustness of the hard-core structure of the soliton.
@@
We also present the nucleon-scalar coupling $g_{S}$ and the scalar susceptibility $C_{\rm NS}$ within the NJL-soliton framework, defined as in Refs. Chanfray2024EPJA ; Chanfray2025 :
@@
We observe a notably small nucleon-scalar coupling [^3] Alternatively, if one evolve the soliton only through the scalar sector of the NJL-soliton model ($F_{\pi}({\cal S}),M_{\pi}({\cal S})$), while keeping the vector parameters fixed at their vacuum values $M_{v}({\cal S})=M_{v}(M_{0}),\ g_{v}({\cal S})=g_{v}(M_{0})$], as done in several works, see Refs. CHRISTOV1990 ; Alkofer1991 , leads to a different nucleon scalar coupling, as discussed in  B. constant within the present framework, accompanied by a negative scalar susceptibility. The appearance of a negative susceptibility is consistent within the NJL-Soliton/skyrmion model. However, we anticipate that incorporating confinement effects, for instance within the chiral confining model, will mitigate this behavior, potentially yielding a positive susceptibility more in line with the results reported in Refs. Chanfray2024EPJA ; Chanfray2025 .
@@
Although we adopt a normalization procedure involving the energy density, specifically multiplying by the factor $\chi$, to reproduce the physical nucleon mass at vacuum (as outlined in Section 2.2), we present and compare the resulting EoS both with and without this approach. The EoSs for the various NJL parameter sets, obtained by eliminating the radial coordinate $r$ from the profiles of energy density $\varepsilon_{\rm slt}(r)$ and pressure $p_{\rm slt}(r)$, are shown in Fig. 7(a), alongside the reference EoS corresponding to the exact parameters of Ref. Fukushima2020 . The impact of the scaling parameter $\chi$ on the EoS is illustrated in Fig. 7(b), and the scaling approach changes the stiffness of the EoSs.
3.2 Effects of Chiral Symmetry Restoration
3.2.1 Scalar field dependency